# Scalable Disk-Based Approximate Nearest Neighbor Search with Page-Aligned Graph

### v1

@@ 1. Introduction
To address this scalability challenge in large-scale vector search, disk-based graph ANNS solutions have been developed (Subramanya et al., 2019; Ni et al., 2023; Pan et al., 2023; Perron et al., 2020; Chen et al., 2021; Guo and Lu, 2025a). These methods offload the data to disk and load it into memory only when needed, thereby reducing memory consumption. For example, the pioneer work DiskANN (Subramanya et al., 2019) store both the full vectors and the graph index on SSDs while retaining compressed vector representations in memory. During search, only the original vectors and the neighbor information along the search path are loaded from disk into memory. Subsequent studies such as SPANN  (Chen et al., 2021), Starting  (Perron et al., 2020), and PipeANN  (Guo and Lu, 2025a) adopt similar designs to further improve query efficiency for large-scale ANN search. However, these methods still face significant I/O latency, particularly under constrained host memory. Moreover, as datasets continue to scale, such approaches encounter persistent scalability bottlenecks.
Figure 1. Impact of memory ratio on query latency for state-of-the-art disk-based ANN schemes (i.e., DiskANN (Subramanya et al., 2019), SPANN (Chen et al., 2021), Starling (Perron et al., 2020), and PipeANN (Guo and Lu, 2025b)) on the SIFT dataset. The memory ratio represents the proportion of the dataset size that is available in host memory. (Due to the minimum memory requirement, PipeANN and SPANN cannot run with the memory ratio below 30%.)
@@ 2.2. Graph-based ANNS
Although graph-based ANNS methods achieve excellent speed and accuracy, they often incur scalability issue due to substantial memory overhead. To address this issue, a range of disk-based graph ANNS solutions has been proposed (Subramanya et al., 2019; Ni et al., 2023; Pan et al., 2023; Perron et al., 2020; Chen et al., 2021; Guo and Lu, 2025a), which reduce memory consumption by offloading data to disk while striving to maintain high search performance. For example, DiskANN(Subramanya et al., 2019), Starling(Perron et al., 2020), and PipeANN(Guo and Lu, 2025b) store original vectors and graph topology on SSDs while retaining only product quantization (PQ)–compressed vectors in memory. In contrast, SPANN(Chen et al., 2021) keeps the graph index entirely in memory and offloads only the vector data to SSDs, performing graph traversal in memory and issuing I/O requests only after traversal completes.
@@ 6.1. Experimental Setup
Baselines. We compare PageANN with four SOTA disk-based ANNS methods including DiskANN (Subramanya et al., 2019), SPANN (Chen et al., 2021), Starling (Perron et al., 2020), and PipeANN (Guo and Lu, 2025b).
@@
Starling (Perron et al., 2020) 1246.94 12.75 157.79 1677.33 9.49 114.53 1008.23 15.80 189.08
PipeANN (Guo and Lu, 2025b) 632.16 25.16 168.51 662.73 14.19 154.47 734.33 23.98 139.52
@@ 7. Related Work
Disk-based ANNS approaches To effectively alleviate the memory constraints of graph-based ANNS, disk-based approaches offload part or all search-related data from memory to disk. A pioneer work, DiskANN (Subramanya et al., 2019), retains only compressed vector values in memory while storing the original vectors and graph index on SSDs. Building on this, DiskANN++ (Ni et al., 2023) and Starling (Perron et al., 2020) optimize SSD page layouts and integrate an in-memory routing process. LM-DiskANN (Pan et al., 2023) offloads all compressed values to disk at the cost of significantly increased disk usage, while PipeANN (Guo and Lu, 2025b) aligns the search algorithm with SSD characteristics. In contrast, SPANN (Chen et al., 2021) keeps the graph in memory and offloads vector values to SSDs, performing graph traversal entirely in memory and issuing all I/O requests only after traversal completes. However, as discussed earlier, all these methods still suffer from high search latency and scalability limitations. PageANN addresses the scalability issue while greatly reducing search latency and increasing throughput. Another types of disk-based ANN schemes focus on the dynamic ANN system to mainly minimize the insertion and graph construction overhead, which is out of the scope with this paper (Xu et al., 2023, 2025; Pound et al., 2025; Zhong et al., 2025).

### v2

@@ 1. Introduction
To address this scalability challenge in large-scale vector search, disk-based graph ANNS solutions have been developed (Subramanya et al., 2019; Ni et al., 2023; Pan et al., 2023; Wang et al., 2024; Chen et al., 2021; Guo and Lu, 2025). These methods offload the data to disk and load it into memory only when needed, thereby reducing memory consumption. For example, the pioneer work DiskANN (Subramanya et al., 2019) store both the full vectors and the graph index on SSDs while retaining compressed vector representations in memory. During search, only the original vectors and the neighbor information along the search path are loaded from disk into memory. Subsequent studies such as SPANN  (Chen et al., 2021), Starting  (Wang et al., 2024), and PipeANN  (Guo and Lu, 2025) adopt similar designs to further improve query efficiency for large-scale ANN search. However, these methods still face significant I/O latency, particularly under constrained host memory. Moreover, as datasets continue to scale, such approaches encounter persistent scalability bottlenecks.
Figure 1. Impact of memory ratio on query latency for state-of-the-art disk-based ANN schemes (i.e., DiskANN (Subramanya et al., 2019), SPANN (Chen et al., 2021), Starling (Wang et al., 2024), and PipeANN (Guo and Lu, 2025)) on the SIFT dataset. The memory ratio represents the proportion of the dataset size that is available in host memory. (Due to the minimum memory requirement, PipeANN and SPANN cannot run with the memory ratio below 30%.)
@@ 2.2. Graph-based ANNS
Although graph-based ANNS methods achieve excellent speed and accuracy, they often incur scalability issue due to substantial memory overhead. To address this issue, a range of disk-based graph ANNS solutions has been proposed (Subramanya et al., 2019; Ni et al., 2023; Pan et al., 2023; Wang et al., 2024; Chen et al., 2021; Guo and Lu, 2025), which reduce memory consumption by offloading data to disk while striving to maintain high search performance. For example, DiskANN(Subramanya et al., 2019), Starling(Wang et al., 2024), and PipeANN(Guo and Lu, 2025) store original vectors and graph topology on SSDs while retaining only product quantization (PQ)–compressed vectors in memory. In contrast, SPANN(Chen et al., 2021) keeps the graph index entirely in memory and offloads only the vector data to SSDs, performing graph traversal in memory and issuing I/O requests only after traversal completes.
@@ 6.1. Experimental Setup
Baselines. We compare PageANN with four SOTA disk-based ANNS methods including DiskANN (Subramanya et al., 2019), SPANN (Chen et al., 2021), Starling (Wang et al., 2024), and PipeANN (Guo and Lu, 2025).
@@
Starling (Wang et al., 2024) 1246.94 12.75 157.79 1677.33 9.49 114.53 1008.23 15.80 189.08
PipeANN (Guo and Lu, 2025) 632.16 25.16 168.51 662.73 14.19 154.47 734.33 23.98 139.52
@@ 7. Related Work
Disk-based ANNS approaches To effectively alleviate the memory constraints of graph-based ANNS, disk-based approaches offload part or all search-related data from memory to disk. A pioneer work, DiskANN (Subramanya et al., 2019), retains only compressed vector values in memory while storing the original vectors and graph index on SSDs. Building on this, DiskANN++ (Ni et al., 2023) and Starling (Wang et al., 2024) optimize SSD page layouts and integrate an in-memory routing process. LM-DiskANN (Pan et al., 2023) offloads all compressed values to disk at the cost of significantly increased disk usage, while PipeANN (Guo and Lu, 2025) aligns the search algorithm with SSD characteristics. In contrast, SPANN (Chen et al., 2021) keeps the graph in memory and offloads vector values to SSDs, performing graph traversal entirely in memory and issuing all I/O requests only after traversal completes. However, as discussed earlier, all these methods still suffer from high search latency and scalability limitations. PageANN addresses the scalability issue while greatly reducing search latency and increasing throughput. Another types of disk-based ANN schemes focus on the dynamic ANN system to mainly minimize the insertion and graph construction overhead, which is out of the scope with this paper (Xu et al., 2023, 2025; Pound et al., 2025; Zhong et al., 2025).